# Forecasting with Hyper-Trees

### v1

@@ Abstract
This paper introduces the concept of Hyper-Trees and offers a new direction in applying tree-based models to time series data. Unlike conventional applications of decision trees that forecast time series directly, Hyper-Trees are designed to learn the parameters of a target time series model. Our framework leverages the gradient-based nature of boosted trees, which allows us to extend the concept of Hyper-Networks to Hyper-Trees and to induce a time-series inductive bias to tree models. By relating the parameters of a target time series model to features, Hyper-Trees address the challenge of parameter non-stationarity and enable tree-based forecasts to extend beyond their initial training range. With our research, we aim to explore the effectiveness of Hyper-Trees across various forecasting scenarios and to expand the application of gradient boosted decision trees past their conventional use in time series forecasting.
@@ 1 Introduction
Contributing to these developments, we propose a novel approach that adapts tree-based models to a time series context via the use of Hyper-Trees. [^2] Even though it shares the same name, our concept of Hyper-Trees is not related to graph theory. Unlike traditional tree-based approaches that model and forecast time series directly, Hyper-Trees focus on learning parameters of target time series models. By reformulating GBDTs into a Hyper-Tree architecture, we combine the benefits of tree-based modeling with the inductive bias of a wide variety of time-series models, thereby extending the capabilities of GBDTs to effectively address the characteristics of time series data. [^3] In its most general form, an inductive bias refers to the set of assumptions a learning algorithm inherently brings to the task of modeling data and which ultimately guides the learning process. In a time series context, these biases are particularly important as they shape how temporal dynamics are modeled and forecasted. For example, a model with an inductive bias towards smooth, gradual changes would be predisposed to forecast time series data with smooth patterns, potentially ignoring abrupt changes or anomalies. As another example, auto-regressive models have a bias towards assuming that future values in or across series can be modeled as a combination of past values. The model choice therefore depends on the nature of the time series data and the specific characteristics one expects to forecast. In contrast to approaches that fit linear models in the terminal nodes of the tree, e.g., the linear-tree approaches of Shi et al., (2019) and de Vito, (2017), Hyper-Trees are more flexible as they allow to parameterize any time series model, without the need to change the code base of existing GBDTs implementations. To the best of our knowledge, we are among the first to transfer the concept of hyper-models to GBDTs. In the remainder of this section, we detail the advantages of using Hyper-Trees for time series modeling and forecasting.
@@
Global Adaptivity of Local Models In the evolving landscape of time series modeling, Hyper-Trees have the potential to redefine the efficiency and adaptability of local forecasting models. Classical local methods like ARIMA (Box et al.,, 2015) or Exponential Smoothing (Holt,, 1957; Winters,, 1960) typically require individual model training for each time series. For instance, in forecasting sales data across various regions, conventional practices would necessitate a distinct model for each region, a process that, while ensuring specificity, does not leverage the cross-series information. Hyper-Trees, in contrast, combine global learning with local adaptability. Rather than treating each time series in isolation, they learn the parameters of a target model across all series. In the sales data scenario, a Hyper-Tree would model all regions collectively, identifying common patterns and trends, and then provide tailored model parameters for each local model individually. Hence, Hyper-Trees allow for parameter sharing across different series, thereby enhancing stability and reliability in model training and forecasting, while also ensuring the local adaptivity of forecasts for each series.
@@
Scalability and Few-shot Learning of Local Models Learning parameters of a univariate local model via Hyper-Trees is particularly useful when individual training of models for each time series becomes impractical due to the volume of the data. Hyper-Trees are capable of learning from a comprehensive dataset and then forecast model parameters for series not originally included in the train set. This cross-learning approach also allows for effective forecasting even in situations with limited historical data, akin to Few-shot Learning (Oreshkin et al.,, 2021; Wang et al.,, 2020; Iwata and Kumagai,, 2020). Such an application of Hyper-Trees is also beneficial in situations where data history is minimal, e.g., forecasting for new geographic locations or for time series with short histories. Here, the Hyper-Tree is trained on time series from existing geographic locations that have sufficient historical data to provide more accurate parameter estimates for the local model, potentially enhancing forecasting reliability for the new series. This flexibility and adaptability of the Hyper-Tree can be particularly advantageous in situations where reliable forecasts are needed from local models, despite the challenges posed by limited historical data availability.
@@
Domain Adaption of Local Models The primary goal of domain adaptation is to develop a model trained within a source domain that can effectively adapt and perform well in a distinct, target domain. This field has predominantly been anchored in deep learning methodologies (Jin et al.,, 2022). Our framework, however, expands this concept to encompass local univariate models through the use of Hyper-Trees. Hyper-Trees play a pivotal role in domain adaptation by dynamically generating or adjusting model parameters. This process utilizes data samples from the target domain as inputs, thereby producing the parameters for a target-specific model. Hence, Hyper-Trees effectively bridge a crucial gap, enabling simpler, more specialized models such as ARIMA (Box et al.,, 2015) or Exponential Smoothing (Holt,, 1957; Winters,, 1960), to harness the sophisticated techniques usually exclusive to deep learning frameworks.
@@
Parameter Estimation Hyper-Trees bring a dynamic approach to parameter estimation, diverging from the static nature of model parameters. Unlike conventional training of time series models, where a set of parameters is estimated that remains fixed for the test set, Hyper-Trees generate parameters at an observational level. For each time step in the test set, they forecast the most appropriate parameters, allowing for adaptability as new information becomes available. This dynamic parameterization enables Hyper-Trees to efficiently respond to a non-stationarity environment commonly present in time series data, without the need of model re-training (Lee et al.,, 2023). Hyper-Trees also offer an advantage in model training by utilizing a more comprehensive approach to parameter estimation. Unlike conventional optimization that mostly relies on first order approximation of the loss function, Hyper-Trees incorporate both gradients and the Hessian, i.e., the curvature of the loss function, enabling more efficient parameter updates.
The key contributions of our approach can be summarized as follows:
@@
  • Expanding the Scope of Tree-Based Models: By transferring the concept of hyper-networks to tree-based models, Hyper-Trees expand the potential applications of tree-based models beyond conventional forecasting tasks. This approach opens new avenues for research and application in the field of time series analysis using a blend of tree-based and time series models.
@@
  • Efficiency in Parameter Space: Hyper-Trees estimate the parameters of a target model using both gradients and the Hessian, potentially leading to more effective forecasting models.
Before we continue with introducing and evaluating Hyper-Trees in the next sections, we would like to stress that the main motivation of this study is not to claim the superiority of our approach over all existing methods. Instead, we aim to initiate an exchange about the potential of tree-based models that goes beyond their traditional use in time series analysis and forecasting. We hope that the ideas and concepts presented here will inspire future research and development in this field and that our exploration of Hyper-Trees serves as a starting point for this. We leave it to the reader to decide whether Hyper-Trees enhance tree-based models by integrating inductive biases from classical time series models, or if they instead broaden classical time series models by incorporating the advantages of tree-based models, encouraging the community to further explore and define their position in the time series domain.
@@ 2.2 Hyper-Tree Architecture
The architecture in Figure 2 shows a unified and modular framework, with the Hyper-Tree as the central component. The Hyper-Tree takes a set of input features and generates parameters $\hat{\theta}_{{}_{\mathcal{H}}}$ that are used by the target model. The target model then processes its input features, e.g., lagged target values, and creates an output $\hat{\psi}_{{}_{\mathcal{T}}}$. This output is subsequently used as input in a shared loss function. The loss function, guided by both the Hyper-Tree and the target model, integrates their outputs to derive gradients and Hessians. This ensures that the loss calculation captures the dynamics of both the Hyper-Tree structure and the target model’s behavior, reflecting the interplay between both components. Contrary to time series models that learn directly from data, the target model in this framework does not. Instead, its role is to induce a specific inductive bias as well as temporal dependencies to the Hyper-Tree. Via gradients and Hessians that are derived from the shared loss function, this bias guides the Hyper-Tree to capture and model the temporal dependencies present in the data by generating parameters $\hat{\theta}_{{}_{\mathcal{H}}}$ that are specific to each time step, effectively encoding the temporal information into the parameters it generates. Consequently, in a Hyper-Tree setup, the target model acts more as a function applying the parameters generated by the Hyper-Tree, rather than learning and internalizing the temporal dependencies on its own. Based on the gradients and Hessians that are derived from the shared loss-function, as well as from the input features it receives for each time step, the Hyper-Tree is the component that learns to encode these dependencies into the target model parameters. Hence, Figure 2 showcases how the target model influences the Hyper-Tree’s learning process, ensuring that the temporal aspects of the data are accurately represented. By reformulating GBDTs into a Hyper-Tree architecture, we combine the benefits of tree-based training of the target-model parameters with the inductive bias of a wide variety of time-series models, thereby extending GBDTs to a time-series context more naturally.
@@
Besides the architectural design shown in Figure 2, we can easily extend the functionality of our Hyper-Tree approach. A common approach to overcome the limitations of tree-based forecasting is to first remove seasonality and trend from the series (temporal part) and then fit a GBDT to the residuals (non-temporal part). The final forecast is then composed of two parts, one that forecasts trend and seasonality, e.g., ARIMA (Box et al.,, 2015), and a tree-based forecast that models the non-temporal part of the series. The rationale behind this decomposition is that it allows for a more precise and targeted modeling of different components of the time series. By decomposing the series into different signals, each model can specialize in capturing the nuances of its respective component, where the non-temporal part includes external feature information, which do not follow regular time patterns but significantly influence the series. However, this approach is not without its drawbacks. Since it involves training two distinct sets of models, one for the temporal and another for the non-temporal part, maintaining two separate models can be complex. More importantly, the independence of each model implies there is no information sharing between the two. This lack of integration implies that potential inter-dependencies between temporal and non-temporal factors are not captured, which might compromise overall forecasting accuracy. Instead of separating the series and training two distinct sets of models, Figure 3 shows that our approach allows for a combination of target model and tree-based forecasts within a unified framework, where $\hat{\psi}_{{}_{\mathcal{T}}}$ models the temporal part of the data, e.g., via an AR-process, $\hat{\psi}_{{}_{\mathcal{H}}}$ models the non-temporal part and $\pi_{{}_{\mathcal{H}}}\in[0,1]$ determines the importance of each component.
@@
In contrast to the two-step approach that is commonly applied, the architecture described in Figure 3 trains a single model within a coherent framework that integrates both temporal and non-temporal elements. Our unified approach enhances the model’s ability to capture interactions between temporal and non-temporal factors, potentially leading to more accurate forecasts.
The modularity of our approach further allows the framework to be extended to a probabilistic setting, where instead of $\hat{\psi}_{{}_{\mathcal{T}}}$ being a point-forecast, the Hyper-Tree architecture can be used to model and forecast parameters of a distribution, from which forecast intervals and quantiles of interest can be derived. [^7] For recent extensions of tree-based models to a probabilistic setting, we refer the interested reader to März and Kneib, (2022); März, (2022, 2019); Sprangers et al., (2021); Hasson et al., (2021); Duan et al., (2020). In general, any suitable distribution that characterizes the data can be specified. As an example, one can assume a Normal distribution $y_{t}\sim N(\mu_{t},\sigma_{t})$ for a HyperTree AR(p), where the mean and standard deviation are parameterized as follows
@@
In this example, we parameterize $\mu_{t}$ via an AR(p) model. Even though we can in principle impose structural assumptions for $\sigma_{t}$ as well, e.g., ARCH/GARCH models (Engle,, 1982; Bollerslev,, 1986), one can, e.g., also directly estimate $\sigma_{t}$ via the Hyper-Tree as illustrated in Figure 4.888Instead of modeling parameters of a distribution, creating probabilistic forecasts via conformal predictive distributions (Vovk et al.,, 2022; Johansson et al.,, 2023) presents an interesting alternative.
@@ 4.1 Local Time Series Models: Illustrative Example
where the parameters $\Theta_{{}_{\mathcal{H}}}(\mathbf{x}_{{}_{t}})=\bigl{[}\theta_{{}_{0,t}}(\mathbf{x}_{{}_{t}}),\theta_{{}_{1,t}}(\mathbf{x}_{{}_{t}}),\theta_{\text{sine}_{i,t}}(\mathbf{x}_{{}_{t}}),\theta_{\text{cosine}_{i,t}}(\mathbf{x}_{{}_{t}})\bigr{]}$ are modeled as functions of time-derived features $\mathbf{x}_{{}_{t}}=\{month,quarter,year,time\}$, with $time$ being a linearly increasing integer. Even though the trend part in Equation (7) is specified as a linear function of $t$, the interactions and non-linearities induced by the Hyper-Tree allow for modeling also complex non-linear patterns. To ensure the trend is a smooth function of time, we add a term that penalizes squared first and second order differences, thereby encouraging neighboring estimates to be close to each other. We represent the seasonality via periodic Fourier-terms that are composed of multiple sine and cosine components, each with different frequencies and amplitudes, where $\theta_{\text{sine}_{i,t}}(\mathbf{x}_{{}_{t}})$ and $\theta_{\text{cosine}_{i,t}}(\mathbf{x}_{{}_{t}})$ represent the sine and cosine weights for the $i$-th seasonal component. For the STL-Hyper-Tree, we set $\text{N}_{\text{season}}=1$ and $p=12$ since the data is recorded on a monthly frequency. For estimating parameters of the target models, we use LightGBM (Ke et al.,, 2017) as a Hyper-Tree model throughout all experiments in this manuscript. [^12] Conceptually, it is straightforward to replace LightGBM (Ke et al.,, 2017) with XGBoost (Chen and Guestrin,, 2016). We opted for LightGBM (Ke et al.,, 2017) as our Hyper-Tree model due to its effectiveness and popularity in the time series community. Instead of tuning the hyper-parameters, we mostly use the default values and train the model for 500 iterations using the Mean-Squared-Error (MSE) as the loss function with the learning rate set to $\eta=0.1$. Figure 5(b) shows the results for the training dataset. As a reference, we add the components of a conventional STL decomposition as available in Seabold and Perktold, (2010).
@@
$\ell_{t}$ $=\alpha_{{}_{t}}(\mathbf{x}_{{}_{t}})\left(y_{t}/s_{t-m}\right)+\left(1-\alpha_{{}_{t}}(\mathbf{x}_{{}_{t}})\right)\left(\ell_{t-1}+\phi_{{}_{t}}(\mathbf{x}_{{}_{t}})b_{t-1}\right)$
$b_{t}$ $=\beta_{{}_{t}}(\mathbf{x}_{{}_{t}})(\ell_{t}-\ell_{t-1})+(1-\beta_{{}_{t}}(\mathbf{x}_{{}_{t}}))\phi_{{}_{t}}(\mathbf{x}_{{}_{t}})b_{t-1}$
$s_{t}$ $=\gamma_{{}_{t}}(\mathbf{x}_{{}_{t}})\frac{y_{t}}{(\ell_{t-1}+\phi_{{}_{t}}(\mathbf{x}_{{}_{t}})b_{t-1})}+(1-\gamma_{{}_{t}}(\mathbf{x}_{{}_{t}}))s_{t-m}$
@@
$\hat{y}_{t+h|t}$ $=\left[\ell_{t}+(\phi_{{}_{t+1}}(\mathbf{x}_{{}_{t}})+\phi^{2}_{{}_{t+2}}(\mathbf{x}_{{}_{t}})+\dots+\phi^{h}_{{}_{t+h}}(\mathbf{x}_{{}_{t}}))b_{t}\right]s_{t+h-m(k+1)}.$ (10)
where $\{\ell_{t},b_{t},s_{t}\}$ denote $\{level,trend,seasonality\}$ with corresponding time-step and feature specific smoothing parameters $\{\alpha(\mathbf{x}_{{}_{t}}),\beta(\mathbf{x}_{{}_{t}}),\gamma(\mathbf{x}_{{}_{t}})\}$, and $m$ indicates the frequency of the seasonality. The choice of the target AR(12) model is not based on any automated model selection. Instead, the decision to use 12 lags stems from the monthly nature of the dataset. Similarly, the trend and seasonality of the airpassenger data have lead us to specify a HyperTree-ES with multiplicative seasonality and damped trend, since this specification often provides accurate and robust forecasts for seasonal data (Hyndman and Athanasopoulos,, 2021). [^15] Even though it is not a Hyper-Network in a strict sense, the Exponential Smoothing Recurrent Neural Network (ESRNN) of Smyl, (2020) can be seen as being related to our HyperTree-ES model. In both implementations, the exponential smoothing part in Equation (10) allows us to capture seasonality, trend, and level, while the LSTM / LightGBM part introduces non-linearities as well as interactions among the parameters. While the HyperTree-ES architecture currently only supports local model training, we plan to extend it to a global setting, enabling cross-learning in order to leverage information at both series and dataset levels. As references, we compare our Hyper-Trees to the following models: [^16] We plan to extend the comparison with other models in later versions of the manuscript. We opted to evaluate the Hyper-Tree architectures to their classical analogs first before we extend the class of models.

### v2

@@ Abstract
This paper introduces the concept of Hyper-Trees and offers a new direction in applying tree-based models to time series data. Unlike conventional applications of decision trees that forecast time series directly, Hyper-Trees are designed to learn the parameters of a target time series model. Our framework leverages the gradient-based nature of boosted trees, which allows us to extend the concept of Hyper-Networks to Hyper-Trees and to induce a time-series inductive bias to tree models. By relating the parameters of a target time series model to features, Hyper-Trees address the issue of parameter non-stationarity and enable tree-based forecasts to extend beyond their training range. With our research, we aim to explore the effectiveness of Hyper-Trees across various forecasting scenarios and to extend the application of gradient boosted decision trees outside their conventional use in time series modeling.
@@ 1 Introduction
Contributing to these developments, we propose a novel approach that adapts tree-based models to a time series context via the use of Hyper-Trees. [^2] Even though it shares the same name, our concept of Hyper-Trees is not related to graph theory. Unlike conventional tree-based approaches that model and forecast time series directly, Hyper-Trees focus on learning parameters of target time series models. By reformulating GBDTs into a Hyper-Tree architecture, we combine the benefits of tree-based modeling with the inductive bias of a wide variety of time series models, thereby extending the capabilities of GBDTs to effectively address the characteristics of time series data. [^3] In its most general form, an inductive bias refers to the set of assumptions a learning algorithm inherently brings to the task of modeling data and which ultimately guides the learning process. In a time series context, these biases are particularly important as they shape how temporal dynamics are modeled and forecasted. For example, a model with an inductive bias towards smooth, gradual changes would be predisposed to forecast time series data with smooth patterns, potentially ignoring abrupt changes or anomalies. As another example, auto-regressive models have a bias towards assuming that future values in or across series can be modeled as a combination of past values. The model choice therefore depends on the nature of the time series data and the specific characteristics one expects to forecast. In contrast to approaches that fit linear models in the terminal nodes of the tree, e.g., the linear-tree approaches of Shi et al., (2019) and de Vito, (2017), Hyper-Trees are more flexible as they allow to parameterize any time series model, without the need to change the code base of existing GBDTs implementations.
To the best of our knowledge, we are among the first to transfer the concept of hyper-models to GBDTs. In the following, we detail the advantages of using Hyper-Trees for time series modeling and forecasting:
@@
  • Global Adaptivity of Local Models In the evolving landscape of time series modeling, Hyper-Trees have the potential to redefine the adaptability of local forecasting models. Classical local methods like ARIMA (Box et al.,, 2015) or Exponential Smoothing (Holt,, 1957; Winters,, 1960) typically require individual model training for each time series. For instance, in forecasting sales data across various regions, conventional practices would necessitate a distinct model for each region, a process that, while ensuring specificity, does not leverage the cross-series information. Hyper-Trees, in contrast, combine global learning with local adaptability. Instead of treating each time series in isolation, they learn the parameters of a target model across all series. In the sales data scenario, a Hyper-Tree would model all regions collectively, identifying common patterns and trends. It then provides tailored model parameters for each region-specific model. Hence, Hyper-Trees allow for parameter sharing across different series, enhancing the stability and reliability of model training and forecasting. At the same time, they maintain the local adaptivity necessary for accurate forecasts for each individual series.
@@
  • Scalability and Few-shot Learning of Local Models Learning parameters of a univariate local model via Hyper-Trees is especially useful when training individual models for each time series becomes impractical due to the large volume of data. Hyper-Trees can learn from a comprehensive dataset and then generate model parameters for time series not originally included in the training set. This cross-learning capability allows for effective forecasting even with limited historical data, similar to Few-shot Learning (Oreshkin et al.,, 2021; Wang et al.,, 2020; Iwata and Kumagai,, 2020). For instance, in scenarios with minimal historical data, such as forecasting for new geographic locations, Hyper-Trees can offer notable benefits. In such a scenario, the Hyper-Tree is trained on time series from existing geographic locations with sufficient historical data, enabling it to generate accurate parameter estimates for the local model applied to the new series. This adaptability of the Hyper-Tree approach ensures more reliable forecasts despite the challenges of limited historical data availability.
@@
  • Domain Adaptation of Local Models The primary goal of domain adaptation is to train a model on data from a source-domain and then adapt it to perform well on data from a related, target-domain. This field has predominantly been anchored in deep learning (Jin et al.,, 2022). Our framework, however, extends this concept to local univariate models. Hyper-Trees play a pivotal role in domain adaptation by dynamically generating or adjusting model parameters based on data from the target-domain. Specifically, a Hyper-Tree trained on data from the source-domain takes data from the target-domain as input and generates the parameters for the target-domain model. This approach allows for the adaptation of simpler models, such as ARIMA (Box et al.,, 2015) or Exponential Smoothing (Holt,, 1957; Winters,, 1960), to different domains by leveraging techniques typically associated with deep learning.
@@
  • Parameter Estimation Hyper-Trees bring a dynamic approach to parameter estimation, diverging from the static nature of model parameters. Unlike conventional training of time series models, where a set of parameters is estimated that remains fixed for the test set, Hyper-Trees generate parameters at an observational level. For each time step in the test set, they forecast the most appropriate parameters, allowing for adaptability as new information becomes available. This dynamic parameterization enables Hyper-Trees to efficiently respond to a non-stationarity environment commonly present in time series data, without the need of model re-training (Lee et al.,, 2023). Hyper-Trees also offer an advantage in model training by utilizing a more comprehensive approach to parameter estimation. Unlike conventional optimization that mostly relies on first order approximation of the loss function, Hyper-Trees incorporate both gradients and the Hessian, i.e., the curvature of the loss function, enabling more efficient parameter updates. In addition, Hyper-Trees offer the advantage over linear parameter estimation of modeling non-linear relationships and interactions between features, enabling them to capture complex patterns in the data. Their built-in regularization further improves parameter estimation and reduces the risk of overfitting. Moreover, Hyper-Trees make it conceptually easy to impose structural assumptions on estimated parameters that allows modeling of, e.g., strictly increasing or decreasing time series patterns. Such assumptions can be imposed by leveraging GBDTs’ monotonicity constraints or by manually adding restrictions on the estimated parameters. This ensures that the estimated parameters adhere to predefined behaviors and align forecasts with expected real-world patterns.
In summary, the key contributions of our Hyper-Tree approach are:
@@
  • Expanding the Scope of Tree-Based Models: By transferring the concept of hyper-networks to tree-based models, Hyper-Trees expand the potential applications of tree-based models beyond their conventional way of forecasting. Our approach opens new avenues for research and application in the field of time series analysis, using a blend of tree-based and classical time series models.
@@
  • Efficiency in Parameter Estimation: Hyper-Trees leverage the strengths of tree-based models to estimate the parameters of a target model, potentially resulting in more effective forecasting models.
Before we continue with introducing and evaluating Hyper-Trees in the next sections, we would like to stress that the main motivation of this study is not to claim the supremacy of our framework over existing models. Rather, our aim is to initiate an exchange about the potential of tree-based models that extends beyond their traditional use in time series analysis and forecasting. We hope that the ideas and concepts presented here will inspire future research and development in this field and that our exploration of Hyper-Trees serves as a starting point for this. We leave it to the reader to decide whether Hyper-Trees enhance tree-based models by integrating inductive biases from classical time series models, or if they instead broaden classical time series models by incorporating the advantages of tree-based models, encouraging the community to further explore and define their position in the time series domain.
@@ 2.2 Hyper-Tree Architecture
The architecture in Figure 2 shows a unified and modular framework, with the Hyper-Tree as the central component. The Hyper-Tree takes a set of input features and generates parameters $\hat{\theta}_{{}_{\mathcal{H}}}$ that are used by the target model. The target model then processes its input features, e.g., lagged target values, and creates an output $\hat{\psi}_{{}_{\mathcal{T}}}$. This output is subsequently used as input in a shared loss function. The loss function, guided by both the Hyper-Tree and the target model, integrates their outputs to derive gradients and Hessians. This ensures that the loss calculation captures the dynamics of both the Hyper-Tree structure and the target model’s behavior, reflecting the interplay between both components. Contrary to time series models that learn directly from data, the target model in this framework does not. Instead, its role is to induce a specific inductive bias as well as temporal dependencies to the Hyper-Tree. Via gradients and Hessians that are derived from the shared loss function, this bias guides the Hyper-Tree to capture and model the temporal dependencies present in the data by generating parameters $\hat{\theta}_{{}_{\mathcal{H}}}$ that are specific to each time step, effectively encoding the temporal information into the parameters it generates. Consequently, in a Hyper-Tree setup, the target model acts more as a function applying the parameters generated by the Hyper-Tree, rather than learning and internalizing the temporal dependencies on its own. Based on the gradients and Hessians that are derived from the shared loss-function, as well as from the input features it receives for each time step, the Hyper-Tree is the component that learns to encode these dependencies into the target model parameters. Figure 2 illustrates how the target model influences the learning process of the Hyper-Tree, ensuring that the temporal aspects of the data are accurately represented. By reformulating well-established GBDTs into a Hyper-Tree architecture, we combine the benefits of tree-based training of the target-model parameters with the inductive bias of a wide variety of time-series models, thereby extending GBDTs to a time-series context more naturally.
@@
Besides the architectural design shown in Figure 2, we can easily extend the functionality of our Hyper-Tree approach. A common approach to overcome the limitations of tree-based forecasting is to first remove seasonality and trend from the series (temporal part) and then fit a GBDT to the residuals (non-temporal part). The final forecast is then composed of two parts, one that forecasts trend and seasonality, e.g., ARIMA (Box et al.,, 2015), and a tree-based forecast that models the non-temporal part of the series. The rationale behind this decomposition is that it allows for a more precise and targeted modeling of different components of the time series. By decomposing the series into different signals, each model can specialize in capturing the nuances of its respective component, where the non-temporal part includes external feature information, which do not follow regular time patterns but significantly influence the series. However, this approach is not without its drawbacks. Since it involves training two distinct sets of models, the independence of each model implies there is no information sharing between the two. This lack of integration implies that potential inter-dependencies between temporal and non-temporal factors are not captured, which might compromise overall forecasting accuracy. Instead of separating the series and training two distinct sets of models, Figure 3 shows that our approach allows for a combination of target model and tree-based forecasts within a unified framework, where $\hat{\psi}_{{}_{\mathcal{T}}}$ models the temporal part of the data, e.g., via an AR-process, $\hat{\psi}_{{}_{\mathcal{H}}}$ models the non-temporal part and $\pi_{{}_{\mathcal{H}}}\in[0,1]$ determines the importance of each component.
@@
In contrast to the two-step approach that is commonly applied, the architecture described in Figure 3 trains a single model within a coherent framework that integrates both temporal and non-temporal elements. Our unified approach enhances the model’s ability to capture interactions between temporal and non-temporal factors, potentially leading to more accurate forecasts. The modularity of our approach further allows the framework to be extended to a probabilistic setting, where instead of $\hat{\psi}_{{}_{\mathcal{T}}}$ being a point-forecast, the Hyper-Tree architecture can be used to model and forecast parameters of a distribution, from which forecast intervals and quantiles of interest can be derived. [^7] For recent extensions of tree-based models to a probabilistic setting, we refer the interested reader to März and Kneib, (2022); März, (2022, 2019); Sprangers et al., (2021); Hasson et al., (2021); Duan et al., (2020). In general, any suitable distribution that characterizes the data can be specified. As an example, one can assume a Normal distribution $y_{t}\sim N(\mu_{t},\sigma_{t})$ for a HyperTree AR($p$), where the mean and standard deviation are parameterized as follows
@@
In this example, we parameterize $\mu_{t}$ via an AR($p$) model. Even though we can in principle impose structural assumptions for $\sigma_{t}$ as well, e.g., ARCH/GARCH models (Engle,, 1982; Bollerslev,, 1986), one can also directly estimate $\sigma_{t}$ via the Hyper-Tree as illustrated in Figure 4.888Instead of modeling parameters of a distribution, creating probabilistic forecasts via conformal predictive distributions (Vovk et al.,, 2022; Johansson et al.,, 2023) presents an interesting alternative.
@@ 4.1 Local Time Series Models: Illustrative Example
where the parameters $\Theta_{{}_{\mathcal{H}}}(\mathbf{x}_{{}_{t}})=\bigl{[}\theta_{{}_{0,t}}(\mathbf{x}_{{}_{t}}),\theta_{{}_{1,t}}(\mathbf{x}_{{}_{t}}),\theta_{\text{sine}_{i,t}}(\mathbf{x}_{{}_{t}}),\theta_{\text{cosine}_{i,t}}(\mathbf{x}_{{}_{t}})\bigr{]}$ are modeled as functions of time-derived features $\mathbf{x}_{{}_{t}}=\{month,quarter,year,time\}$, with $time$ being a linearly increasing integer. Even though the trend part in Equation (7) is specified as a linear function of $t$, the interactions and non-linearities induced by the Hyper-Tree allow for modeling also complex non-linear patterns. To ensure the trend is a smooth function of time, we add a term that penalizes squared first and second order differences, thereby encouraging neighboring estimates to be close to each other. We represent the seasonality via periodic Fourier-terms that are composed of multiple sine and cosine components, each with different frequencies and amplitudes, where $\theta_{\text{sine}_{i,t}}(\mathbf{x}_{{}_{t}})$ and $\theta_{\text{cosine}_{i,t}}(\mathbf{x}_{{}_{t}})$ represent the sine and cosine weights for the $i$-th seasonal component. For the STL-Hyper-Tree, we set $\text{N}_{\text{season}}=1$ and $p=12$ since the data is recorded on a monthly frequency. For estimating parameters of the target models, we use LightGBM (Ke et al.,, 2017) as a Hyper-Tree model throughout all experiments in this manuscript. [^12] Conceptually, it is straightforward to replace LightGBM (Ke et al.,, 2017) with XGBoost (Chen and Guestrin,, 2016). We opted for LightGBM (Ke et al.,, 2017) as our Hyper-Tree model due to its effectiveness and popularity in the time series community. Instead of tuning the hyper-parameters, we use the default values and train the model for 500 iterations using the Mean-Squared-Error (MSE) as the loss function with the learning rate set to $\eta=0.1$. Figure 5(b) shows the results for the training dataset. As a reference, we add the components of a conventional STL decomposition as available in Seabold and Perktold, (2010).
@@
$\ell_{t}$ $=\alpha_{{}_{t}}(\mathbf{x}_{{}_{t}})\left(y_{t}/s_{t-m}\right)+\bigl{(}1-\alpha_{{}_{t}}(\mathbf{x}_{{}_{t}})\bigr{)}\bigl{(}\ell_{t-1}+\phi_{{}_{t}}(\mathbf{x}_{{}_{t}})b_{t-1}\bigr{)}$
$b_{t}$ $=\beta_{{}_{t}}(\mathbf{x}_{{}_{t}})(\ell_{t}-\ell_{t-1})+\bigl{(}1-\beta_{{}_{t}}(\mathbf{x}_{{}_{t}})\bigr{)}\phi_{{}_{t}}(\mathbf{x}_{{}_{t}})b_{t-1}$
$s_{t}$ $=\gamma_{{}_{t}}(\mathbf{x}_{{}_{t}})\frac{y_{t}}{\bigl{(}\ell_{t-1}+\phi_{{}_{t}}(\mathbf{x}_{{}_{t}})b_{t-1}\bigr{)}}+\bigl{(}1-\gamma_{{}_{t}}(\mathbf{x}_{{}_{t}})\bigr{)}s_{t-m}$
@@
$\hat{y}_{t+h|t}$ $=\left[\ell_{t}+\bigl{(}\phi_{{}_{t+1}}(\mathbf{x}_{{}_{t+1}})+\phi^{2}_{{}_{t+2}}(\mathbf{x}_{{}_{t+2}})+\dots+\phi^{h}_{{}_{t+h}}(\mathbf{x}_{{}_{t+h}})\bigr{)}b_{t}\right]s_{t+h-m(k+1)}.$ (10)
where $\{\ell_{t},b_{t},s_{t}\}$ denote $\{level,trend,seasonality\}$ with corresponding time-step and feature specific smoothing parameters $\{\alpha(\mathbf{x}_{{}_{t}}),\beta(\mathbf{x}_{{}_{t}}),\gamma(\mathbf{x}_{{}_{t}})\}$, $m$ indicates the frequency of the seasonality and $\{\phi_{{}_{t+1}}(\mathbf{x}_{{}_{t+1}}),\ldots,\phi_{{}_{t+h}}(\mathbf{x}_{{}_{t+h}})\}$ control the dampening of the trend. The choice of the target AR(12) model is not based on any automated model selection. Instead, the decision to use 12 lags stems from the monthly nature of the dataset. Similarly, the trend and seasonality of the air-passenger dataset have lead us to specify a HyperTree-ES with multiplicative seasonality and damped trend, since this specification often provides accurate and robust forecasts for seasonal data (Hyndman and Athanasopoulos,, 2021). [^15] Even though it is not a Hyper-Network in a strict sense, the Exponential Smoothing Recurrent Neural Network (ESRNN) of Smyl, (2020) can be seen as being related to our HyperTree-ES model. In both implementations, the exponential smoothing part in Equation (10) allows us to capture seasonality, trend, and level, while the LSTM / LightGBM part introduces non-linearities as well as interactions among the parameters. While the HyperTree-ES architecture currently only supports local model training, we plan to extend it to a global setting, enabling cross-learning in order to leverage information at both series and dataset levels. As references, we compare our Hyper-Trees to the following models: [^16] We plan to extend the comparison with other models in later versions of the manuscript. We opted to evaluate the Hyper-Tree architectures to their classical analogs first before we extend the class of models.
@@
  • LightGBM-AR(12): same as LightGBM but with additional autoregressive lagged-target features.